ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 35, 76 Socolowsky, Jürgen, Fachhochschule Brandenburg, Germany On a perturbed two-layer free surface flow down an inclined plane Mathematical models of multiple-layer coating flows with viscous liquids usually represent free boundary value problems for the stationary Navier-Stokes equations in unbounded two-dimensional domains. In many cases the flow domains have two outlets to infinity. Using methods of nonlinear analysis the solvability of a free boundary value problem for the flow down an inclined plane with a small perturbation can be shown. Part of these results is joint work with K. Pileckas. Reference: K. Pileckas, and J. Socolowsky, Analysis of the NavierStokes equations for some two-layer flows in unbounded domains, in: Proceedings of the 6th International Conference on Navier-Stokes Equations and related Nonlinear Problems, eds. H. Amann et al., VSP/TEV (1998), (accepted for publication)